OpenStudy (anonymous):

I need help on a precalc question, medal and fan: Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 6. Please just walk me through it so I know how to do it.

OpenStudy (anonymous):

@ganeshie8 can you get someone on here to help me?

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

please dude

OpenStudy (phi):

do you know the generic equation for an ellipse?

OpenStudy (anonymous):

x^2/a^2 + y^2/b^2 = 1

OpenStudy (phi):

ok, and generally the a is the semi major-diameter (1/2 of the longer diameter) if the y direction is longer, you put the a^2 underneath the y

OpenStudy (anonymous):

ok so now what values do I have to plug in?

OpenStudy (phi):

vertical major axis of length 18, that is telling you the y direction is the major axis the 18 is the entire height of the ellipse. you want 1/2 of that for "a"

OpenStudy (phi):

minor axis of length 6 that is the width of the ellipse. you want 1/2 of that for b

OpenStudy (anonymous):

so 9^2 =81 y^2/81 + x^2/b^2 = 1

OpenStudy (phi):

so far, so good

OpenStudy (anonymous):

so now what about b?

OpenStudy (phi):

I posted about "b" up above.

OpenStudy (anonymous):

oh lol sorry, so 3^2 = 9 y^2/81 + x^2/9 = 1?

OpenStudy (phi):

yes

OpenStudy (anonymous):

can you help with another?

OpenStudy (anonymous):

@ phi ? Find all polar coordinates of point P where P = (5,-pi/6)